what is 2log5x - log5 2= log5(2x+6)
Answer by dabanfield(803)

what is 2log5x - log5 2= log5(2x+6)
2log5 x = log5 x^2 so the left side above becomes:
log5 x^2 - log5 2 = log5 (x^2/2)
We have then:
log5 (x^2/2) = log5 (2x+6) so
x^2/2 = 2x+6
x^2 = 2*(2x+6)
x^2 = 4x + 12
x^2 - 4x - 12 = 0
(x-6)*(x+2) = 0
x-6 = 0 or x+2 = 0
x = 6 or x = -2